OpenStudy (anonymous):

PLEASE HELP!!!!!!! Josh believes the Spanish club students at his school have an unfair advantage in being assigned to the Spanish class they request. He asked 500 students at his school the following questions: "Are you in the Spanish club?" and "Did you get the Spanish class you requested?" The results are shown in the table below: Help Josh determine if all students at his school have an equal opportunity to get the Spanish class they

OpenStudy (anonymous):

Spanish Club Not in Spanish Club Total Received Spanish class requested 265 100 365 Did not get Spanish class requested 70 65 135 Total 335 165 500

OpenStudy (anonymous):

Help Josh determine if all students at his school have an equal opportunity to get the Spanish class they requested. Show your work, and explain your process for determining the fairness of the class assignment process.

OpenStudy (kropot72):

You can calculate the following probabilities: 1. Given that a sampled student is in the Spanish Club, what is the probability they got the Spanish class they requested? 2. Given that a sampled student is not in the Spanish Club, what is the probability they got the Spanish class they requested? If there is a significant difference between the two probabilities, it indicates there is a bias in the selection procedure.

OpenStudy (anonymous):

So, would it be p(Spanish class they requested)= 265/365? would it be p(not in the Spanish Club)= 100/365?

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

Given that a sampled student is in the Spanish Club, the probability they got the Spanish class they requested is given by 265/335. Given that a sampled student is not in the Spanish Club, the probability they got the Spanish class they requested is given by 100/165.

OpenStudy (anonymous):

Was 365 the wrong value? Why 335 and 165?

OpenStudy (kropot72):

335 is the number of sampled students who are in the Spanish club. 165 is the number of sampled students who are not in the Spanish club. We are looking at only the first two columns in the table.

OpenStudy (anonymous):

Oh, ok. I get, had to re-read the question

OpenStudy (anonymous):

now the question asked: Help Josh determine if all students at his school have an equal opportunity to get the Spanish class they requested.

OpenStudy (anonymous):

so would I just use the first one column??

OpenStudy (kropot72):

If a student is in the Spanish club, the probability they got the Spanish class they requested is 265/335 = 0.79. If a student is not in the Spanish club, the probability they got the Spanish class they requested is 100/165 = 0.61. Do you think this result is fair?

OpenStudy (anonymous):

no, based on that it's not.

OpenStudy (anonymous):

So, would that be the solution?

OpenStudy (kropot72):

I am satisfied that the reasoning answers the question. All students do not have an equal chance of getting into the Spanish class that they requested.

OpenStudy (anonymous):

I can't put that lol but I'll have to explain it in more depth but thank you. :) Can I ask another question?

OpenStudy (kropot72):

You're welcome. I suggest that you use my first post as a basis for your explanation. Please post your other question as a new question.